OpenStudy (anonymous):

Sum or difference (H). Let a and b be any two irrational numbers. Show that either a+b or a-b must be irrational

OpenStudy (anonymous):

i think you can do this by assuming that \(a+b\) is rational, say \(a+b=r\) and then showing that \(a-b\) is irrational

OpenStudy (anonymous):

or maybe easier by contradiction. suppose both are rational say if \(a+b=r_2\) and \(a-b=r_2\) where \(r_1,r_2\) are rational then \[a+b+a-b=2a=r_1+r_2\] so \(2a\) is rational as the sum of two rational numbers is also rational therefore \(a=\frac{r_1+r_2}{2}\) is also rational contradicting the assumption that \(a\) is irrational